14. Draw the structure of the MAJOR product(s) formed in each of the following reactions.
(i) 3-methylpentane + Br2 + light →
(ii) 2-methylhexan-3-ol + H2SO4 →
(iii) cyclopentyl ethyl ether + excess HBr
(iv) hexan-3-one + excess CH3OH/H2SO4 →
(v) octanol + pyridinium chlorochromate →
21

CHE1502 OCT NOV 2020
(vi) pentyne + H2O/H2SO4/HgSO4 →
(vii) 3-cyclohexylbutene + HBr
(viii) 1-methylcyclopentene + BH3/THF → followed by addition of H2O2/NaOH
(ix) 4,4-dimethylpent-1-ene + Hg(OAc)2/H2O → followed by the addition of NaBH4
(x) n-heptylamine + HCl →
1
Expert's answer
2020-12-08T05:33:37-0500

14.

(i)CH3-CH2-CH2 (CH3) - CH2-CH3+Br2=CH3-CH2-CHBr (CH3) - CH2-CH3+HBr bromination to produce 3-bromo-3-methylpentane

(ii)СH3-CH(CH3)-CH(OH)-CH2-CH2-CH3+H2SO4=СH3-CH(CH3)-CH(SO4)-CH2-CH2-CH3+H20

(iiI)




(iV)


(v)



21

(vi)



(vii)




(viii)





(ix)




(x)